OpenStudy (anonymous):

I ran into a 1st order DiffEQ that....

OpenStudy (anonymous):

I cannot figure which method to solve it with (linear, exact, homo, ect) x(dy/dx)=y + sqrt(x^2 + y^2)

OpenStudy (anonymous):

\[x \frac{ dy }{ dx }=y + \sqrt( x^2-y^2) \]

OpenStudy (anonymous):

subtraction not addition, sorry...

OpenStudy (mimi_x3):

homogenous \(y=ux\) you have to change it; to seperable

OpenStudy (mimi_x3):

\[\large x\frac{dy}{dx}=y+\sqrt{x^2+y^2}\] correct?

OpenStudy (anonymous):

no x^2-y^2 is under the radical, sorry about typo

OpenStudy (mimi_x3):

\[\large x\frac{dy}{dx}=y+\sqrt{x^2-y^2}\]???

OpenStudy (anonymous):

yes.

OpenStudy (mimi_x3):

so, its homogeous \(y=ux\) \[\frac{dy}{dx}=v+x\frac{dv}{dx}\] are you able to start it?

OpenStudy (anonymous):

indeed i have it from here.... thank you for the help!

OpenStudy (mimi_x3):

:) if you need any help; feel free to ask :)